# A Low-Rank ADMM Splitting Approach for Semidefinite Programming

### 14 Mar 2024 v1

@@ A.4 Proof of Lemma 1
$\begin{split}\|C^{k+1}-C^{k}\|\|\lambda^{k}\|\leq&\frac{1}{\eta}\big{(}L_{f}\delta\sqrt{s_{A}}+2\gamma\delta\sqrt{s_{A}}\big{)}\|\Delta U^{k+1}\|\end{split}$ (39)
@@
$\begin{split}\eta\|{\lambda}^{k+1}-{\lambda}^{k}\|\leq&\big{(}L\delta^{2}+L_{f}+\gamma+\frac{1}{\eta}\big{(}L_{f}\delta\sqrt{s_{A}}+2\gamma\delta\sqrt{s_{A}}\big{)}\big{)}\|\Delta U^{k+1}\|\\ &+\big{(}L\delta^{2}+\gamma\big{)}\|\Delta V^{k+1}\|.\end{split}$
@@
$\begin{split}\|{\lambda}^{k+1}-{\lambda}^{k}\|^{2}\leq&\frac{2}{\eta^{2}}\big{(}L\delta^{2}+L_{f}+\gamma+\frac{1}{\eta}L_{f}\delta\sqrt{s_{A}}+\frac{2}{\eta}\gamma\eta\sqrt{s_{A}}\big{)}^{2}\|\Delta U^{k+1}\|^{2}\\ &+\frac{2}{\eta^{2}}\big{(}L\delta^{2}+\gamma\big{)}^{2}\|\Delta V^{k+1}\|^{2}\\ \triangleq&c_{1}\|\Delta U^{k+1}\|^{2}+c_{2}\|\Delta V^{k+1}\|^{2}\end{split}$ (40)
@@
$,\begin{split}&\mathcal{L}_{\rho}(U^{k+1},V^{k+1},{\lambda}^{k+1})-\mathcal{L}_{\rho}(U^{k},V^{k},{\lambda}^{k})\\ \leq&\frac{1}{\rho}\|\Delta{\lambda}^{k+1}\|^{2}-\frac{\gamma}{2}\big{(}\|\Delta U^{k+1}\|^{2}+\|\Delta V^{k+1}\|^{2}\big{)}\\ \leq&-(\frac{\gamma}{2}-\frac{c_{1}}{\rho})\|\Delta U^{k+1}\|^{2}-(\frac{\gamma}{2}-\frac{c_{2}}{\rho})\|\Delta V^{k+1}\|^{2}\\ \leq&0.\end{split}$
@@
$\begin{split}L_{\rho}(U^{k+1},V^{k+1},{\lambda}^{k+1})+{1\over\rho}\|\lambda^{k+1}\|^{2}&\leq L_{\rho}(U^{k},V^{k},{\lambda}^{k})+{1\over\rho}\|\lambda^{k+1}\|^{2}\\ &\leq\beta^{0}+{1\over\rho}\left({L_{f}\delta+2\gamma\delta\over\eta}\right)^{2}<\hat{\beta}.\end{split}$
@@ A.5 Proof of Theorem 2
Based on the developments outlined in Lemma 1, we deduce the existence of a constant $\kappa$ such that
$\begin{split}\mathcal{L}_{\rho}(U^{k+1},V^{k+1},{\lambda}^{k+1})-\mathcal{L}_{\rho}(U^{k},V^{k},{\lambda}^{k})\leq-\kappa\|\Delta U^{k+1}\|^{2}-\kappa\|\Delta V^{k+1}\|^{2}.\end{split}$ (41)
@@
$\|\nabla\mathcal{L}(U^{k+1},V^{k+1},\lambda^{k+1})\|_{F}^{2}\leq\left({L_{f}^{\delta}+L^{\delta}\delta^{2}+\rho s_{A}\delta^{2}+\sqrt{s_{A}}\delta}+\gamma\right)^{2}\|\Delta V^{k+1}\|_{F}^{2}+{1\over\rho^{2}}\|\Delta\lambda^{k+1}\|^{2},$
@@
$\mathcal{L}_{\rho}(U,V,\lambda)=\mathcal{L}(U,V,\lambda)+{\rho\over 2}\|\mathcal{A}(UV^{T})-b\|^{2},$
@@
$\|\nabla\mathcal{L}_{\rho}(U^{k+1},V^{k+1},\lambda^{k+1})\|_{F}\leq\zeta(\|\Delta U^{k+1}\|+\|\Delta V^{k+1}\|+\|\Delta\lambda^{k+1}\|).$ (42)
Therefore, by using (42), (41) and Assumption 4, we can invoke [Attouch et al., 2013, Theorem 2.9] that the sequence $\{(U^{k},V^{k},\lambda^{k})\}$ converges to a critical point of $L_{\rho}(U,V,\lambda)$, which corresponds to a KKT point of problem (6). We further have the sequence $\{(U^{k},V^{k},\lambda^{k})\}$ has a finite length, i.e.
$\sum_{j=1}^{\infty}\|\Delta U^{j+1}\|_{F}+\|\Delta V^{j+1}\|_{F}+\|\Delta\lambda^{j+1}\|<\infty.$
@@
$\min_{1\leq j\leq K}\|\Delta U^{j+1}\|_{F}+\|\Delta V^{j+1}\|_{F}+\|\lambda^{j+1}-\lambda^{j}\|=O({1\over K}).$
@@
$\begin{split}\min_{1\leq j\leq K}&\left\{\|\nabla f(U^{k+1}(V^{k+1})^{\top})V^{k+1}+\gamma(U^{k+1}-V^{k+1})+\sum_{i=1}^{m}\lambda_{i}^{k+1}A_{i}V^{k+1}\|_{F}\right.\\ &\,+\|\nabla f(V^{k+1}(U^{k+1})^{\top})V^{k+1}+\gamma(V^{k+1}-U^{k+1})+\sum_{i=1}^{m}\lambda_{i}^{k+1}A_{i}V^{k+1}\|_{F}\\ &\,+\|\mathcal{A}(U^{k+1}(V^{k+1})^{T})-b\|\bigg{\}}=O\left(\frac{1}{k}\right).\end{split}$
@@ A.6 Proof of Theorem 3
Because of the convergence of the sequence $\{(U^{k},V^{k},\lambda^{k})\}_{k=0}^{\infty}$, the iteration point will enter the local area around the limit point and satisfies the Łojasiewicz inequality after a constant number of steps. For the simplicity of notation, we assume $(U^{k},V^{k},\lambda^{k})\in\mathcal{B}_{\delta_{c}}(\overline{U},\overline{V},\overline{\lambda})$ for any $k\geq 0$ in the following analysis.
@@
$\phi^{\prime}(\mathcal{L}_{\rho}(U,V,{\lambda})-\mathcal{L}_{\rho}(\overline{U},\overline{V},\overline{\lambda}))\cdot\|\nabla\mathcal{L}_{\rho}(U,V,{\lambda})\|\geq 1.$ (43)
@@
$\phi^{\prime}(l^{k})\cdot\|\nabla\mathcal{L}_{\rho}(U^{k},V^{k},{\lambda}^{k})\|\geq 1.$
@@
$\phi^{\prime}(l^{k})(l^{k}-l^{k+1})\cdot\|\nabla\mathcal{L}_{\rho}(U^{k},V^{k},{\lambda}^{k})\|\geq l^{k}-l^{k+1}.$
@@
$\big{(}\phi(l^{k})-\phi(l^{k+1})\big{)}\cdot\|\nabla\mathcal{L}_{\rho}(U^{k},V^{k},{\lambda}^{k})\|\geq l^{k}-l^{k+1}.$
According to Lemma 1 and (42), we can bound $l^{k}-l^{k+1}$ and $\|\nabla\mathcal{L}_{\rho}(U^{k},V^{k},{\lambda}^{k})\|$ respectively
$l^{k}-l^{k+1}\geq\kappa\big{(}\|\Delta U^{k+1}\|^{2}+\|\Delta V^{k+1}\|^{2}\big{)},$
$\|\nabla\mathcal{L}_{\rho}(U^{k},V^{k},{\lambda}^{k})\|\leq\zeta\big{(}\|\Delta U^{k}\|+\|\Delta V^{k}\|+\|\Delta\lambda^{k}\|\big{)}.$
Note that $\|\Delta{\lambda}^{k+1}\|^{2}$ can be upper bounded by $\|\Delta U^{k+1}\|^{2}$ and $\|\Delta V^{k+1}\|^{2}$, hence there exists some constant $\tilde{\kappa}>0$ such that
$l^{k}-l^{k+1}\geq\tilde{\kappa}\big{(}\|\Delta U^{k+1}\|^{2}+\|\Delta V^{k+1}\|^{2}+\|\Delta\lambda^{k+1}\|^{2}\big{)}.$
@@
$\frac{\zeta}{\tilde{\kappa}}\big{(}\phi(l^{k})-\phi(l^{k+1})\big{)}\big{(}\|\Delta U^{k}\|+\|\Delta V^{k}\|+\|\Delta\lambda^{k}\|\big{)}\geq\big{(}\|\Delta U^{k+1}\|^{2}+\|\Delta V^{k+1}\|^{2}+\|\Delta\lambda^{k+1}\|^{2}\big{)}.$
@@
$\frac{2\zeta}{\tilde{\kappa}}\big{(}\phi(l^{k})-\phi(l^{k+1})\big{)}+\big{(}\|\Delta U^{k}\|+\|\Delta V^{k}\|+\|\Delta\lambda^{k}\|\big{)}\geq 2\big{(}\|\Delta U^{k+1}\|+\|\Delta V^{k+1}\|+\|\Delta\lambda^{k+1}\|\big{)}.$
@@
$\frac{2\zeta}{\tilde{\kappa}}\big{(}\phi(l^{k})-\phi(l^{K+1})\big{)}+\big{(}\|\Delta U^{k}\|+\|\Delta V^{k}\|+\|\Delta\lambda^{k}\|\big{)}\geq$ $\sum_{t=k}^{K}\big{(}\|\Delta U^{t+1}\|+\|\Delta V^{t+1}\|+\|\Delta\lambda^{t+1}\|\big{)}$
$+\big{(}\|\Delta U^{K+1}\|+\|\Delta V^{K+1}\|+\|\Delta\lambda^{K+1}\|\big{)}.$
@@
$\frac{2\zeta}{\tilde{\kappa}}\phi(l^{k})+\big{(}\|\Delta U^{k}\|+\|\Delta V^{k}\|+\|\Delta\lambda^{k}\|\big{)}\geq\sum_{t=k}^{\infty}\big{(}\|\Delta U^{t+1}\|+\|\Delta V^{t+1}\|+\|\Delta\lambda^{t+1}\|\big{)}.$ (44)
Let $S^{k}=\sum_{t=k}^{\infty}\big{(}\|\Delta U^{t+1}\|+\|\Delta V^{t+1}\|+\|\Delta\lambda^{t+1}\|\big{)}$. Then we can rewrite the above inequality as follows
@@
$c(l^{k})^{-\alpha}\cdot\|\nabla\mathcal{L}_{\rho}(U^{k},V^{k},{\lambda}^{k})\|\geq 1,$
@@
$\phi(l^{k})=\frac{c}{1-\alpha}(l^{k})^{1-\alpha}\leq\tilde{c}\big{(}\|\Delta U^{k}\|+\|\Delta V^{k}\|+\|\Delta\lambda^{k}\|\big{)}^{\frac{1-\alpha}{\alpha}}=\tilde{c}(S^{k-1}-S^{k})^{\frac{1-\alpha}{\alpha}}.$
@@ Appendix B Additional Numerical Results
In this section, we provide the solving errors for the instances tested in Section 5 in the following tables.

### 25 Mar 2024 v2

@@ A.4 Proof of Lemma 1
$\begin{split}\|C^{k+1}-C^{k}\|_{F}\|\lambda^{k}\|_{2}\leq&\frac{1}{\eta}\big{(}L_{f}\delta\sqrt{s_{A}}+2\gamma\delta\sqrt{s_{A}}\big{)}\|\Delta U^{k+1}\|_{F}\end{split}$ (39)
@@
$\begin{split}\eta\|{\lambda}^{k+1}-{\lambda}^{k}\|_{2}\leq&\big{(}L\delta^{2}+L_{f}+\gamma+\frac{1}{\eta}\big{(}L_{f}\delta\sqrt{s_{A}}+2\gamma\delta\sqrt{s_{A}}\big{)}\big{)}\|\Delta U^{k+1}\|_{F}\\ &+\big{(}L\delta^{2}+\gamma\big{)}\|\Delta V^{k+1}\|_{F}.\end{split}$
@@
$\begin{split}\|{\lambda}^{k+1}-{\lambda}^{k}\|_{2}^{2}\leq&\frac{2}{\eta^{2}}\big{(}L\delta^{2}+L_{f}+\gamma+\frac{1}{\eta}L_{f}\delta\sqrt{s_{A}}+\frac{2}{\eta}\gamma\eta\sqrt{s_{A}}\big{)}^{2}\|\Delta U^{k+1}\|_{F}^{2}\\ &+\frac{2}{\eta^{2}}\big{(}L\delta^{2}+\gamma\big{)}^{2}\|\Delta V^{k+1}\|_{F}^{2}\\ \triangleq&c_{1}\|\Delta U^{k+1}\|_{F}^{2}+c_{2}\|\Delta V^{k+1}\|_{F}^{2}\end{split}$ (40)
@@
$,\begin{split}&\mathcal{L}_{\rho}(U^{k+1},V^{k+1},{\lambda}^{k+1})-\mathcal{L}_{\rho}(U^{k},V^{k},{\lambda}^{k})\\ \leq&\frac{1}{\rho}\|\Delta{\lambda}^{k+1}\|_{2}^{2}-\frac{\gamma}{2}\big{(}\|\Delta U^{k+1}\|_{F}^{2}+\|\Delta V^{k+1}\|^{2}\big{)}\\ \leq&-(\frac{\gamma}{2}-\frac{c_{1}}{\rho})\|\Delta U^{k+1}\|_{F}^{2}-(\frac{\gamma}{2}-\frac{c_{2}}{\rho})\|\Delta V^{k+1}\|_{F}^{2}\\ \leq&0.\end{split}$
@@
$\begin{split}L_{\rho}(U^{k+1},V^{k+1},{\lambda}^{k+1})+{1\over\rho}\|\lambda^{k+1}\|_{2}^{2}&\leq L_{\rho}(U^{k},V^{k},{\lambda}^{k})+{1\over\rho}\|\lambda^{k+1}\|_{2}^{2}\\ &\leq\beta^{0}+{1\over\rho}\left({L_{f}\delta+2\gamma\delta\over\eta}\right)^{2}<\hat{\beta}.\end{split}$
@@ A.5 Proof of Theorem 2
Proof Based on the developments outlined in Lemma 1, we deduce the existence of a constant $\kappa$ such that
$\begin{split}\mathcal{L}_{\rho}(U^{k+1},V^{k+1},{\lambda}^{k+1})-\mathcal{L}_{\rho}(U^{k},V^{k},{\lambda}^{k})\leq-\kappa\|\Delta U^{k+1}\|_{F}^{2}-\kappa\|\Delta V^{k+1}\|_{F}^{2}.\end{split}$ (41)
@@
$\|\nabla\mathcal{L}(U^{k+1},V^{k+1},\lambda^{k+1})\|_{F}^{2}\leq\left({L_{f}^{\delta}+L^{\delta}\delta^{2}+\rho s_{A}\delta^{2}+\|\mathcal{A}\|_{2}\delta}+\gamma\right)^{2}\|\Delta V^{k+1}\|_{F}^{2}+{1\over\rho^{2}}\|\Delta\lambda^{k+1}\|^{2},$
@@
$\mathcal{L}_{\rho}(U,V,\lambda)=\mathcal{L}(U,V,\lambda)+{\rho\over 2}\|\mathcal{A}(UV^{T})-b\|_{2}^{2},$
@@
$\|\nabla\mathcal{L}_{\rho}(U^{k+1},V^{k+1},\lambda^{k+1})\|_{F}\leq\zeta(\|\Delta U^{k+1}\|_{F}+\|\Delta V^{k+1}\|_{F}+\|\Delta\lambda^{k+1}\|_{2}).$ (42)
Therefore, by using (42), (41) and Assumption 4, we can invoke [Attouch et al., 2013, Theorem 2.9] that the sequence $\{(U^{k},V^{k},\lambda^{k})\}_{k=0}^{\infty}$ converges to a critical point of $L_{\rho}(U,V,\lambda)$, which corresponds to a KKT point of problem (6). We further have the sequence $\{(U^{k},V^{k},\lambda^{k})\}_{k=0}^{\infty}$ has a finite length, i.e.
$\sum_{j=1}^{\infty}\|\Delta U^{j+1}\|_{F}+\|\Delta V^{j+1}\|_{F}+\|\Delta\lambda^{j+1}\|_{2}<\infty.$
@@
$\min_{1\leq j\leq K}\|\Delta U^{j+1}\|_{F}+\|\Delta V^{j+1}\|_{F}+\|\Delta\lambda^{j+1}\|_{2}=O({1\over K}).$
@@
$\begin{split}\min_{1\leq j\leq K}&\left\{\|\nabla f(U^{k+1}(V^{k+1})^{\top})V^{k+1}+\gamma(U^{k+1}-V^{k+1})+\sum_{i=1}^{m}\lambda_{i}^{k+1}A_{i}V^{k+1}\|_{F}\right.\\ &\,+\|\nabla f(V^{k+1}(U^{k+1})^{\top})V^{k+1}+\gamma(V^{k+1}-U^{k+1})+\sum_{i=1}^{m}\lambda_{i}^{k+1}A_{i}V^{k+1}\|_{F}\\ &\,+\|\mathcal{A}(U^{k+1}(V^{k+1})^{T})-b\|_{2}\bigg{\}}=O\left(\frac{1}{K}\right).\end{split}$
@@ A.6 Proof of Theorem 3
Proof Because of the convergence of the sequence $\{(U^{k},V^{k},\lambda^{k})\}_{k=0}^{\infty}$, the iteration point will enter the local area around the limit point and satisfies the Łojasiewicz inequality after a constant number of steps. For the simplicity of notation, we assume $(U^{k},V^{k},\lambda^{k})\in\mathcal{B}_{\delta_{c}}(\overline{U},\overline{V},\overline{\lambda})$ for any $k\geq 0$ in the following analysis.
@@
$\phi^{\prime}(\mathcal{L}_{\rho}(U,V,{\lambda})-\mathcal{L}_{\rho}(\overline{U},\overline{V},\overline{\lambda}))\cdot\|\nabla\mathcal{L}_{\rho}(U,V,{\lambda})\|_{F}\geq 1.$ (43)
@@
$\phi^{\prime}(l^{k})\cdot\|\nabla\mathcal{L}_{\rho}(U^{k},V^{k},{\lambda}^{k})\|_{F}\geq 1.$
@@
$\phi^{\prime}(l^{k})(l^{k}-l^{k+1})\cdot\|\nabla\mathcal{L}_{\rho}(U^{k},V^{k},{\lambda}^{k})\|_{F}\geq l^{k}-l^{k+1}.$
@@
$\big{(}\phi(l^{k})-\phi(l^{k+1})\big{)}\cdot\|\nabla\mathcal{L}_{\rho}(U^{k},V^{k},{\lambda}^{k})\|_{F}\geq l^{k}-l^{k+1}.$
According to Lemma 1 and (42), we can bound $l^{k}-l^{k+1}$ and $\|\nabla\mathcal{L}_{\rho}(U^{k},V^{k},{\lambda}^{k})\|_{F}$ respectively
$l^{k}-l^{k+1}\geq\kappa\big{(}\|\Delta U^{k+1}\|_{F}^{2}+\|\Delta V^{k+1}\|_{F}^{2}\big{)},$
$\|\nabla\mathcal{L}_{\rho}(U^{k},V^{k},{\lambda}^{k})\|_{F}\leq\zeta\big{(}\|\Delta U^{k}\|_{F}+\|\Delta V^{k}\|_{F}+\|\Delta\lambda^{k}\|_{2}\big{)}.$
Note that $\|\Delta{\lambda}^{k+1}\|_{2}^{2}$ can be upper bounded by $\|\Delta U^{k+1}\|_{F}^{2}$ and $\|\Delta V^{k+1}\|_{F}^{2}$, hence there exists some constant $\tilde{\kappa}>0$ such that
$l^{k}-l^{k+1}\geq\tilde{\kappa}\big{(}\|\Delta U^{k+1}\|_{F}^{2}+\|\Delta V^{k+1}\|_{F}^{2}+\|\Delta\lambda^{k+1}\|_{2}^{2}\big{)}.$
@@
$\frac{\zeta}{\tilde{\kappa}}\big{(}\phi(l^{k})-\phi(l^{k+1})\big{)}\big{(}\|\Delta U^{k}\|_{F}+\|\Delta V^{k}\|_{F}+\|\Delta\lambda^{k}\|_{2}\big{)}\geq\big{(}\|\Delta U^{k+1}\|_{F}^{2}+\|\Delta V^{k+1}\|_{F}^{2}+\|\Delta\lambda^{k+1}\|_{2}^{2}\big{)}.$
@@
$\frac{2\zeta}{\tilde{\kappa}}\big{(}\phi(l^{k})-\phi(l^{k+1})\big{)}+\big{(}\|\Delta U^{k}\|_{F}+\|\Delta V^{k}\|_{F}+\|\Delta\lambda^{k}\|_{2}\big{)}\geq 2\big{(}\|\Delta U^{k+1}\|_{F}+\|\Delta V^{k+1}\|_{F}+\|\Delta\lambda^{k+1}\|_{2}\big{)}.$
@@
$\frac{2\zeta}{\tilde{\kappa}}\big{(}\phi(l^{k})-\phi(l^{K+1})\big{)}+\big{(}\|\Delta U^{k}\|_{F}+\|\Delta V^{k}\|_{F}+\|\Delta\lambda^{k}\|_{2}\big{)}\geq$ $\sum_{t=k}^{K}\big{(}\|\Delta U^{t+1}\|_{F}+\|\Delta V^{t+1}\|_{F}+\|\Delta\lambda^{t+1}\|_{2}\big{)}$
$+\big{(}\|\Delta U^{K+1}\|_{F}+\|\Delta V^{K+1}\|_{F}+\|\Delta\lambda^{K+1}\|_{2}\big{)}.$
@@
$\frac{2\zeta}{\tilde{\kappa}}\phi(l^{k})+\big{(}\|\Delta U^{k}\|_{F}+\|\Delta V^{k}\|_{F}+\|\Delta\lambda^{k}\|_{2}\big{)}\geq\sum_{t=k}^{\infty}\big{(}\|\Delta U^{t+1}\|_{F}+\|\Delta V^{t+1}\|_{F}+\|\Delta\lambda^{t+1}\|_{2}\big{)}.$ (44)
Let $S^{k}=\sum_{t=k}^{\infty}\big{(}\|\Delta U^{t+1}\|_{F}+\|\Delta V^{t+1}\|_{F}+\|\Delta\lambda^{t+1}\|_{2}\big{)}$. Then we can rewrite the above inequality as follows
@@
$c(l^{k})^{-\alpha}\cdot\|\nabla\mathcal{L}_{\rho}(U^{k},V^{k},{\lambda}^{k})\|_{F}\geq 1,$
@@
$\phi(l^{k})=\frac{c}{1-\alpha}(l^{k})^{1-\alpha}\leq\tilde{c}\big{(}\|\Delta U^{k}\|_{F}+\|\Delta V^{k}\|_{F}+\|\Delta\lambda^{k}\|_{2}\big{)}^{\frac{1-\alpha}{\alpha}}=\tilde{c}(S^{k-1}-S^{k})^{\frac{1-\alpha}{\alpha}}.$
@@ Appendix B Additional Numerical Results
In this section, we provide the solving errors for the instances tested in Section 5. Since each solver terminates with distinct criteria, we present three commonly defined errors for the solutions of each solver including:
Primal Infeasibility:
$\frac{\|\mathcal{A}(X)-b\|_{2}}{1+\|b\|_{1}},$
Dual Infeasibility:
$\frac{|\min\{0,\sigma_{\min}(C-\mathcal{A}^{*}(\lambda))\}|}{1+\|c\|_{1}},$
Primal Dual Gap:
$\frac{\langle C,X\rangle-\lambda^{\top}b}{1+|\langle C,X\rangle|+|\lambda^{\top}b|}.$
The three error measures are chosen for ranking in Hans Mittelmann’s benchmarks, and it is also widely used in many existing solvers, like COPT [Ge et al., 2022] and HDSDP [Gao et al., 2022]. The reports of errors in all three experiments are shown in the following tables.